OpenStudy (anonymous):

When Maggie applies the brakes of her car, the car slows uniformly from 14.7 m/s to 0 m/s in 2.36 s. How far ahead of a stop sign must she apply her brakes in order to stop at the sign? Answer in units of m

OpenStudy (anonymous):

\[x_f=x_i+v(t_f-t_i)\] \[x_f=0+14.7(2.36-0)\] \[x_f=34.692_m\]